How much is your arbitrage profit in Euros at expiration, if you know that the current exchange rate is Euro 1.2 / GBP, the 60-day forward rate is Euro 1.12 /GBP, the risk free rate in Europe is 2% and the risk free rate in the UK is 4%? Make your calculation so that the spot transaction is for GBP 500,000.

Please provide your arbitrage profit in Euros rounded to two decimals.
Answer #1
Forward rate= EURO per GBP * (1+Interest rate in EURO)/(1+Interest rate in GBP)
Forward rate= =1.20*(1+0.02)/(1+0.04)
Per GBP              1.1769
Since given forward rate is 1.12, we can see that some arbitrage opportunity is available
If fund invested in GBP so amount= =500000*(1+4%)
£   520,000.00
If amount invested in EURO
Equivelent Euro= 500000*1.20
Equivelent Euro=            600,000
The amount payable in EURO =600000*(1+2%)
           612,000
The amount converted back to GBP= 612000/1.12
The amount converted back to GBP= £   546,428.57
So this way there will be gain £     26,428.57